Suppose you want to combine two types of fruit drink to create 24 kg of a drink that will be 5% sugar by weight. Fruit drink A is 4% sugar weight, and fruit drink B is 8%percent sugar by weight
Answer by addingup(3677)

So you are telling me that:
A+B = 24 If I subtract A from 24 I have B. Agree?
B = 24-A We'll use this value for B next.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
0.04A+0.08B = 0.05(24) Now substitute for B:
0.04A+0.08(24-A) = 0.05(24) Multiply on left and right to get rid of parenthesis:
0.04A+1.92-0.08A = 1.2 Subtract A on left and 1.92 on both sides:
-0.04A = -0.72 divide both sides by -0.04 and remember that -/- = +
A = 18 You need 18kg of A and:
24-18 = 6 kg of B
Happy learning.
J